OpenStudy (anonymous):

blinn college in bryan has 620 students. the difference between male students and female students is 20. how many female students are there?

OpenStudy (anonymous):

m-f=20 m=20+f 620=m+f 620=(20+f)+f 620=20+2f 600=2f 300=f